Arrange the following groups from lowest to highest entropy (Assume 1 mole of each at the same conditions.): A)H2O(l), CHCL3(l), CCl4(g),C(s)

B) NaBr(s), NaI(s), KI(s) , LiBr(s)
Answer #1

The molecule which possess more disorderness have highest entropy. Gas molecules have highest entropy than liquid and which higher entropy than solid molecules. Entropy order is solid<liquid<gas.

liquid cloroform has entropy of 201.7J/K greaterthan water

a) C(s) < H2O(l)<CHCl3(l)<CCl4(g)

b)here all molecules are solid. Among those

entropy of NaBr=86.82J/K, NaI=98.5J/K, KI=106.39J/K, LiBr=74.1J/K.

ie; LiBr<NaBr<NaI<KI